# Universality of Efimov states in highly mass-imbalanced cold-atom mixtures with van der Waals and dipole interactions

### v2

@@ III.3 Moderate dipole interaction $a_{dd}\sim r_{\mathrm{vdw}}$
In Fig. 8, we show the binding energies of the fermionic Efimov states plotted against $K_{c,F}^{(3D)}$ as obtained above. The numerical data of different short-range $R_{\mathrm{min}}$ values collapse into a single curve, demonstrating the universality of the fermionic Efimov states. The universality holds particularly well for a realistic $a_{dd}$ value of 167Er-167Er (Fig. 8(a)(b)), where only a few points of higher angular momentum character deviate from the universal curve. In contrast to the bosonic case, we cannot find an analytical expression which could explain the universal energy spectra owing to the impossibility of using $v_{p}^{(\mathrm{HH})}$ hence Eq. (16). The avoided crossing at $K_{c,F}^{(3D)}\simeq-1$ originates from the $\ell=3$ channel which is present without the dipole interaction at around $r_{\mathrm{vdw}}/v_{p}^{(\mathrm{HH})}\simeq-0.3$ (see Fig. 2(c) and Fig. 12 in Appendix D). It shifts toward a larger negative $K_{c,F}^{(3D)}$ side with a decreased slope as the dipole strength is increased from a realistic value of 167Er-167Er (Fig. 8(a)(b)) toward larger values (Fig. 8(c)-(f)). The foothill of the avoided crossing shows a larger variance for stronger dipole interactions (Fig. 8(c)-(f)). This is likely be a precursor of the breakdown of our assumption that the single-channel short-range phase in $\ell=1$ universally characterizes the three-body system (see Sec. III.4), especially for a very strong dipole interaction and large $K_{c,F}^{(3D)}$ (Fig. 8(c)(e)(f)).
@@
Although the $K_{c,F}^{(3D)}$ parameter has the advantage of capturing the universal behavior of the fermionic Efimov states, it is marginally useful in predicting the three-body parameter of experimental systems of interest, e.g. 167Er-167Er-6Li. This is primarily because it is not easy, though not impossible, to determine the value of $K_{c,F}^{(3D)}$; it sensitively depends on the position (i.e. $R_{\mathrm{min}}$) and shape of the short-range repulsive potential, and it is therefore challenging to predict its value by theoretical calculations [57, 65, 66]. Furthermore, as the $K_{c,F}^{(3D)}$ dependence appears as the next-next-next leading order if we perform the asymptotic expansion of the $J_{3}$ and $Y_{3}$ at large distance in Eq. (19), it contributes to the low-energy scattering cross section as a small correction term modifying the dominant quasi-universal contribution [60].
@@ III.4 Strong dipole interaction $a_{dd}\gg r_{\mathrm{vdw}}$
Figure 9: Binding energy of the bosonic (blue circles) and fermionic (orange squares) Efimov states for strong dipole interactions of $a_{dd}/r_{\mathrm{vdw}}=50$ (top row (a)(b)), and $a_{dd}/r_{\mathrm{vdw}}=100$ (bottom row (c)(d)). $K_{c}^{(1D)}$ in the horizontal axis is obtained numerically from Eqs. (20) (21). The mass ratio is taken as $M/m=27.5855...$ which corresponds to 166Er-6Li for the boson, while for the fermion it is hypothetically taken to be the same value as the boson to demonstrate the difference via the statistics of the particles.
@@
In Fig. 9, the binding energies of the Efimov states obtained from Eq. (17) are plotted as a function of $K_{c}^{(1D)}$ in Eqs. (20) (21). The bosonic (blue circles) and fermionic (orange squares) systems, hypothetically assumed to be the same mass ratio, show rather similar behaviors, in stark contrast to the weak and moderate dipole systems of Sec. III.1-III.3. This is because the parity of the wavefunction plays a marginal role when the two heavy atoms are mostly aligned in a quasi-one-dimensional configuration along the $z$-axis, and the $z>0$ and $z<0$ regions are separated by the hard-wall boundary condition. This is particularly true when the size of the trimer is much smaller than $a_{dd}$, beyond which the centrifugal repulsion in Eq. (17) representing the three-dimensional nature becomes more dominant than the dipole term. For the tightly bound states in Fig. 9, the bosonic and fermionic systems become almost indistinguishable, supporting this scenario. In other words, the statistics of the particles is irrelevant for the strongly dipolar Efimov states as long as their binding energy is larger than the dipole scale $1/Ma_{dd}^{2}$.
@@
Figure 10: Contour plot of $|r\psi(\bm{r})|$ of the relative position of the two heavy particles in the Efimov states, plotted in the $y=0$ plane. (a) Shallow bound states in the universal regime for the bosons, and (b) for the fermions. (c) Shallow bound states in the non-universal regime (i.e, a steep slope region in-between the avoided crossings in Fig. 9(a)) for the bosons, and (d) tightly bound states for the bosons (plotted in a smaller $xz$ region than (a)-(c)). For all (a)-(d), the dipole strength and mass ratio are the same as those in Fig. 9 (a)(b), with the following parameters: (a) $Mr_{\mathrm{vdw}}^{2}|E|=1.24\times 10^{-4}$, $(K_{c}^{(1D)})^{-1}=0.71...$ ($R_{\mathrm{min}}/r_{\mathrm{vdw}}=0.25225...$), (b) $Mr_{\mathrm{vdw}}^{2}|E|=1.23\times 10^{-4}$, $(K_{c}^{(1D)})^{-1}=2.58...$ ($R_{\mathrm{min}}/r_{\mathrm{vdw}}=0.2617...$), (c) $Mr_{\mathrm{vdw}}^{2}|E|=1.10\times 10^{-4}$, $(K_{c}^{(1D)})^{-1}=-1.14...$ ($R_{\mathrm{min}}/r_{\mathrm{vdw}}=0.25758...$), and (d) $Mr_{\mathrm{vdw}}^{2}|E|=9.72\times 10^{-3}$, $(K_{c}^{(1D)})^{-1}=-3.42...$ ($R_{\mathrm{min}}/r_{\mathrm{vdw}}=0.2592...$).
@@
The data points with a small slope in Fig. 9, including the weakly bound states, tend to collapse into a single curve for both the bosonic and fermionic systems. Although the spreading of the data points is larger than those in Figs. 5 and 8, it remains much smaller than the Efimov period, even with an almost 16 times change in the short-range part of the van der Waals interaction. We have also confirmed that the spreading gets smaller as $R_{\mathrm{min}}$ is decreased, in a similar manner to Fig. 5. We can thus conclude that $K_{c}^{(1D)}$ universally characterizes the short-range phase, hence the Efimov states with the strong dipole interactions. On the other hand, the data points with steeper slopes do not seem to show the universality with respect to $K_{c}^{(1D)}$. Since those states have significant contributions from large angular momentum states $\ell\gg 1$ and the angular insensitive nature of their short-range phase can break down, it is not unexpected to find that their features cannot be universally captured by the one-dimensional scattering model in Eqs. (20) (21) where the angular momentum is absent (see also dimers in Fig. 13 in Appendix D with similar behaviors). Nevertheless, the physical mechanics behind the universal and non-universal behaviors in Sec. III.3 and Fig. 9 and how they change in-between them have yet to be clarified.
@@
In Fig. 10, we show the amplitude of the wavefunction. For a bosonic Efimov state with a small binding energy (Fig. 10 (a)), it essentially comprises of two regions; the inner region $R\lesssim a_{dd}$ where the dipole interaction is so strong that the two heavy particles are preferentially aligned on the $z$-axis, and the outer region $R\gtrsim a_{dd}$ where the dipole interaction is smaller than the kinetic term, thus forming a three-dimensional halo state of a predominantly $s$-wave nature. They are neither a perfect one-dimensional geometry nor a spherical $s$-wave halo owing to the strong admixture of angular momentum states, but their universal nature is essentially determined by the phase in the short-range region, where the one-dimensional configuration is dominant and hence $K_{c}^{(1D)}$ is a universal parameter. For fermionic Efimov states with a similarly small value of the binding energy (Fig. 10 (b)), the short-range and long-range natures are almost the same, with the only notable difference being the presence of a node in the wavefunction at $z=0$, representing the anti-symmetrization. In contrast, the wavefunction in the non-universal regime (Fig. 10 (c)) shows a distinct feature: the wavefunction in the inner region $R\lesssim a_{dd}$ no longer shows a simple quasi-one-dimensional character, but rather a strongly anisotropic shape, which is likely induced by an admixture of various angular momentum states. This is consistent with the breakdown of the universality with respect to $K_{c}^{(1D)}$. When the binding energy is so large that the size of the wavefunction is smaller than $a_{dd}$ (Fig. 10 (d)), the whole wavefunction is localized in the inner region $R\lesssim a_{dd}$, forming an almost perfect quasi-one-dimensional state. In this limit, the hard-wall boundary condition $R=R_{\mathrm{min}}$ necessarily produces a node in the wavefunction at $z\approx 0$, so that the bososnic and fermionic wavefunctions become the same except for the sign change. In this limit, the bosonic and fermionic systems show the same universal behaviors. While this limit of boson-fermion duality, analogous to the Tonks-Girardeau limit [68, 69], is fascinating, it is challenging to achieve because it requires a very large dipole interaction strength, which is impossible for cold atoms with magnetic dipole interactions, but may be possible with polar molecules [70, 71] or Rydberg atoms [72, 73].
@@ IV Three-body parameters of Er-Er-Li Efimov states
In Table 1, we list some combinations of isotopes of the Er-Li systems, together with their physical parameters. The dipole lengths are comparable to the Er-Er van der Waals length $a_{dd}/r_{\mathrm{vdw}}\simeq 0.87$, which lie outside the perturbative regime $a_{dd}/r_{\mathrm{vdw}}\lesssim 0.5$ discussed in Sec. III.2. Nevertheless, they are not strong enough to be in the strong-dipole limit discussed in Sec. III.4; they are in the parameter region well described by the renormalized van der Waals universality shown in Sec. III.3. We can therefore use Eq. (15) to evaluate the binding energy of the Efimov states for the bosonic Er atoms interacting with a 6Li atom in unitary limit. The $s$-wave scattering length between the Er atoms $a^{(\mathrm{HH})}$ generally depends on the applied magnetic field, owing to the existence of a series of broad and narrow Feshbach resonances [57, 65], and it is yet to be determined by experiments or by sophisticated theoretical calculations. Still, by using the background $s$-wave scattering lengths reported in the literature [57, 58] as representative values, we can estimate the binding energy of the Efimov states $|E|=\kappa_{*}^{2}/M$. The results are tabulated in Table 1; we find that the difference in the values of $a^{(\mathrm{HH})}$ significantly affect the three-body parameter $\kappa_{*}$ across the isotopes while the mass and $a_{dd}$ are almost the same and therefore play marginal roles. The values of $\kappa_{*}$ in Table 1 can be readily improved with better knowledge of the Er-Er scattering length $a^{(\mathrm{HH})}$. We also note that the larger binding energy states in Table 1 may not be so accurate because we use the low-energy analytical formula in Eq. (15); this can also be straightforwardly improved using the numerical results of Sec. III.3.
Table 1: Three-body parameters of the Er-Er-Li Efimov states: binding energy at the unitary Er-Li scattering length $|E|=\kappa_{*}^{2}/M$, and the Er-Li scattering length $a_{-}^{(\mathrm{HL})}$ ($a_{*}^{(\mathrm{HL})}$) at which the Efimov states dissociate into three atoms (an atom and a dimer), respectively. The van der Waals and dipole lengths between the Er atoms are from Refs [57, 58] with the mass scaling. $a^{(\mathrm{HH})}$ between the heavy atoms are estimated by the background scattering lengths $a_{bg}$ tabulated in Refs [58]. $a_{0}$ is the Bohr radius.
@@ V Conclusion & Outlook
We have studied the universality of the Efimov states in a mass-imbalanced three-body system of heavy-heavy-light atoms with a unitary interaction between the heavy-light atoms, with the major motivation of elucidating the interplay between the van der Waals and magnetic dipole interactions in Er-Er-Li Efimov states. By reducing the three-body problem to a coupled-channel two-body equation between the heavy atoms using the Born-Oppenheimer approximation, we have investigated analytically and numerically the three-body parameters (i.e. binding energies in the unitary limit) of the Efimov states for various strengths of dipole interactions between the bosonic and fermionic heavy atoms.
@@
In the absence of dipole interaction, using the quantum defect theory, we derive analytical formulae of the three-body parameters of the Efimov states, which are represented as explicit universal functions of the van der Waals length and $s$-wave scattering length ($p$-wave scattering volume) between the heavy bosonic (fermionic) atoms. The analytical formulae excellently reproduce the numerical results reported in Ref. [25], demonstrating the van der Waals universality in a analytical manner.
@@
In the presence of dipole interaction, different orbital angular momentum states get coupled and we have performed coupled-channel numerical calculations. For a weak dipole interaction, we find that the three-body parameters are affected by the dipole interaction in a different manner for the bosons and fermions. Due to the presence (absence) of the dipole matrix element for bosons in the $L=0$ (fermions in $L=1$) states, the three-body parameter gets a quadratic (linear) shift with respect to a perturbative dipole interaction for the bosons (fermions), respectively.
@@
When the dipole interaction is as strong as the van der Waals interaction, we find that the values of the three-body parameters get significantly altered from those of a non-dipolar system. However, once the $s$-wave scattering length between the heavy atoms is calculated in the presence of both the dipole and van der Waals interactions, the QDT formulae derived for a pure van der Waals system can well reproduce the three-body parameters of the bosonic Efimov states. Using this renormalized van der Waals universality, we have estimated the three-body parameters of the Er-Er-Li Efimov states for the bososnic Er isotopes. For the fermionic heavy atoms, on the other hand, the $p$-wave scattering volume can no longer be defined, owing to the long-range nature of the dipole interaction, so that the renormalized van der Waals universality does not hold. However, they can still be described universally with a low-energy scattering parameter between the two fermionic dipoles which characterizes the phase shift of an asymptotic $p$-wave state, so that the universality of the three-body parameter holds true.
@@
For an extremely large dipole interaction, neither the $s$-wave scattering length nor the $p$-wave low-energy scattering parameter can universally characterize the Efimov states, because the two dipoles get preferentially aligned in a linear geometry at short distance and they are neither $s$-wave nor $p$-wave dominant. In this limit, a quasi-one-dimensional scattering parameter is capable of capturing the short-range scattering phase. We show that except for some regions in-between avoided crossings, the Efimov states can be universally described by the one-dimensional scattering parameter.
@@
With the above studies, we have elucidated how the interplay of the van der Waals and dipole interactions can affect the universality of the Efimov states. In particular, we have clarified the roles of the statistics of the particles and the angular momentum channels, in the presence of both isotropic and anisotropic forces. Our work is not only relevant for the current challenge of observing Efimov states in the highly mass-imbalanced cold atom mixture of Er-Li atoms [41, 42], but also can shed new light on halo nuclear phenomena. The Efimov states are believed to appear universally in various strongly interacting quantum systems [18, 8, 3]. In particular, Efimov-like weakly bound three-body states have been actively studied in nuclear systems, such as neutron-rich nuclei [7, 80, 81, 82], Hoyle state of 12[83, 84], and in near-breakup excited states of nuclei [85]. In nuclear physics, the tensor force has been pointed out to play a crucial role in determining the stability and reactions of nuclei. Notably, the tensor force has an anisotropic nature similar to the dipole-dipole interaction [86]. Investigating the interplay of the van der Waals interaction and dipole interaction is thus analogous to the nuclear problems where the interplay of the central part of the strong force and the tensor force is relevant. Our study on the dipole and van der Waals interactions in strongly interacting quantum few-body systems lays the basis for the quantum simulation of nuclear phenomena using dipolar cold-atom mixtures [41, 42, 43, 58, 87].

### v3

@@ III.3 Moderate dipole interaction $a_{dd}\sim r_{\mathrm{vdw}}$
In Fig. 8, we show the binding energies of the fermionic Efimov states plotted against $K_{c,F}^{(3D)}$ as obtained above. The numerical data of different short-range $R_{\mathrm{min}}$ values collapse into a single curve, demonstrating the universality of the fermionic Efimov states. The universality holds particularly well for a realistic $a_{dd}$ value of 167Er-167Er (Fig. 8(a)(b)), where only a few points of higher angular momentum character deviate from the universal curve. In contrast to the bosonic case, we cannot find an analytical expression which could explain the universal energy spectra owing to the impossibility of using $v_{p}^{(\mathrm{HH})}$ and hence Eq. (16). The avoided crossing at $K_{c,F}^{(3D)}\simeq-1$ originates from the $\ell=3$ channel which is present without the dipole interaction at around $r_{\mathrm{vdw}}^{3}/v_{p}^{(\mathrm{HH})}\simeq-0.3$ (see Fig. 2(c) and Fig. 12 in Appendix D). It shifts toward a larger negative $K_{c,F}^{(3D)}$ side with a decreased slope as the dipole strength is increased from a realistic value of 167Er-167Er (Fig. 8(a)(b)) toward larger values (Fig. 8(c)-(f)). The foothill of the avoided crossing shows a larger variance for stronger dipole interactions (Fig. 8(c)-(f)). This is likely be a precursor of the breakdown of our assumption that the single-channel short-range phase in $\ell=1$ universally characterizes the three-body system (see Sec. III.4), especially for a very strong dipole interaction and large $K_{c,F}^{(3D)}$ (Fig. 8(c)(e)(f)).
@@
Although the $K_{c,F}^{(3D)}$ parameter has the advantage of capturing the universal behavior of the fermionic Efimov states, it is marginally useful in predicting the three-body parameter of experimental systems of interest, e.g., 167Er-167Er-6Li. This is primarily because it is not easy, though not impossible, to determine the value of $K_{c,F}^{(3D)}$; it sensitively depends on the position (i.e. $R_{\mathrm{min}}$) and shape of the short-range repulsive potential, and it is therefore challenging to predict its value by theoretical calculations [57, 65, 66]. Furthermore, as the $K_{c,F}^{(3D)}$ dependence appears as the next-next-next leading order if we perform the asymptotic expansion of the $J_{3}$ and $Y_{3}$ at large distance in Eq. (19), it contributes to the low-energy scattering cross section as a small correction term modifying the dominant quasi-universal contribution [60].
@@ III.4 Strong dipole interaction $a_{dd}\gg r_{\mathrm{vdw}}$
Figure 9: Binding energy of the bosonic (blue circles) and fermionic (orange squares) Efimov states for strong dipole interactions of $a_{dd}/r_{\mathrm{vdw}}=50$ (top row (a)(b)), and $a_{dd}/r_{\mathrm{vdw}}=100$ (bottom row (c)(d)). $K_{c}^{(1D)}$ in the horizontal axis is obtained numerically from Eqs. (20) and (21). The mass ratio is taken as $M/m=27.5855...$, which corresponds to 166Er-6Li for the boson, while for the fermion it is hypothetically taken to be the same value as the boson to demonstrate the difference via the statistics of the particles.
@@
In Fig. 9, the binding energies of the Efimov states obtained from Eq. (17) are plotted as a function of $K_{c}^{(1D)}$ in Eqs. (20) and (21). The bosonic (blue circles) and fermionic (orange squares) systems, hypothetically assumed to be the same mass ratio, show rather similar behaviors, in stark contrast to the weak and moderate dipole systems of Secs. III.1-III.3. This is because the parity of the wave function plays a marginal role when the two heavy atoms are mostly aligned in a quasi-one-dimensional configuration along the $z$-axis, and the $z>0$ and $z<0$ regions are separated by the hard-wall boundary condition. This is particularly true when the size of the trimer is much smaller than $a_{dd}$, beyond which the centrifugal repulsion in Eq. (17) representing the three-dimensional nature becomes more dominant than the dipole term. For the tightly bound states in Fig. 9, the bosonic and fermionic systems become almost indistinguishable, supporting this scenario. In other words, the statistics of the particles is irrelevant for the strongly dipolar Efimov states as long as their binding energy is larger than the dipole scale $1/Ma_{dd}^{2}$.
@@
Figure 10: Contour plot of $|r\psi(\bm{r})|$ of the relative position of the two heavy particles in the Efimov states, plotted in the $y=0$ plane. (a) Shallow bound states in the universal regime for the bosons, and (b) for the fermions. (c) Shallow bound states in the non-universal regime (i.e, a steep slope region between the avoided crossings in Fig. 9(a)) for the bosons, and (d) tightly bound states for the bosons (plotted in a smaller $xz$ region than (a)-(c)). For all (a)-(d), the dipole strength and mass ratio are the same as those in Fig. 9 (a)(b), with the following parameters: (a) $Mr_{\mathrm{vdw}}^{2}|E|=1.24\times 10^{-4}$, $(K_{c}^{(1D)})^{-1}=0.71...$ ($R_{\mathrm{min}}/r_{\mathrm{vdw}}=0.25225...$), (b) $Mr_{\mathrm{vdw}}^{2}|E|=1.23\times 10^{-4}$, $(K_{c}^{(1D)})^{-1}=2.58...$ ($R_{\mathrm{min}}/r_{\mathrm{vdw}}=0.2617...$), (c) $Mr_{\mathrm{vdw}}^{2}|E|=1.10\times 10^{-4}$, $(K_{c}^{(1D)})^{-1}=-1.14...$ ($R_{\mathrm{min}}/r_{\mathrm{vdw}}=0.25758...$), and (d) $Mr_{\mathrm{vdw}}^{2}|E|=9.72\times 10^{-3}$, $(K_{c}^{(1D)})^{-1}=-3.42...$ ($R_{\mathrm{min}}/r_{\mathrm{vdw}}=0.2592...$).
@@
The data points with a small slope in Fig. 9, including the weakly bound states, tend to collapse into a single curve for both the bosonic and fermionic systems. Although the spreading of the data points is larger than those in Figs. 5 and 8, it remains much smaller than the Efimov period, even with an almost 16 times change in the short-range part of the van der Waals interaction. We have also confirmed that the spreading gets smaller as $R_{\mathrm{min}}$ is decreased, in a similar manner to Fig. 5. We can thus conclude that $K_{c}^{(1D)}$ universally characterizes the short-range phase, hence the Efimov states with the strong dipole interactions. On the other hand, the data points with steeper slopes do not seem to show the universality with respect to $K_{c}^{(1D)}$. Since those states have significant contributions from large angular momentum states $\ell\gg 1$ and the angular insensitive nature of their short-range phase can break down, it is not unexpected to find that their features cannot be universally captured by the one-dimensional scattering model in Eqs. (20) and (21) where the angular momentum is absent (see also dimers in Fig. 13 in Appendix D with similar behaviors). Nevertheless, the physical mechanism behind the universal and non-universal behaviors in Sec. III.3 and Fig. 9 and how they change between them have yet to be clarified.
@@
In Fig. 10, we show the amplitude of the wave function. For a bosonic Efimov state with a small binding energy (Fig. 10 (a)), it essentially comprises two regions; the inner region $R\lesssim a_{dd}$ where the dipole interaction is so strong that the two heavy particles are preferentially aligned on the $z$ axis, and the outer region $R\gtrsim a_{dd}$ where the dipole interaction is smaller than the kinetic term, thus forming a three-dimensional halo state of a predominantly $s$-wave nature. They are neither a perfect one-dimensional geometry nor a spherical $s$-wave halo owing to the strong admixture of angular momentum states, but their universal nature is essentially determined by the phase in the short-range region, where the one-dimensional configuration is dominant and hence $K_{c}^{(1D)}$ is a universal parameter. For fermionic Efimov states with a similarly small value of the binding energy (Fig. 10 (b)), the short-range and long-range natures are almost the same, with the only notable difference being the presence of a node in the wave function at $z=0$, representing the anti-symmetrization. In contrast, the wave function in the non-universal regime (Fig. 10 (c)) shows a distinct feature: the wave function in the inner region $R\lesssim a_{dd}$ no longer shows a simple quasi-one-dimensional character, but rather a strongly anisotropic shape, which is likely induced by an admixture of various angular momentum states. This is consistent with the breakdown of the universality with respect to $K_{c}^{(1D)}$. When the binding energy is so large that the size of the wave function is smaller than $a_{dd}$ (Fig. 10 (d)), the whole wave function is localized in the inner region $R\lesssim a_{dd}$, forming an almost perfect quasi-one-dimensional state. In this limit, the hard-wall boundary condition $R=R_{\mathrm{min}}$ necessarily produces a node in the wave function at $z\approx 0$, so that the bososnic and fermionic wave functions become the same except for the sign change. In this limit, the bosonic and fermionic systems show the same universal behaviors. While this limit of boson-fermion duality, analogous to the Tonks-Girardeau limit [68, 69], is fascinating, it is challenging to achieve because it requires a very large dipole interaction strength, which is impossible for cold atoms with magnetic dipole interactions, but may be possible with polar molecules [70, 71] or Rydberg atoms [72, 73].
@@ IV Three-body parameters of Er-Er-Li Efimov states
In Table 1, we list some combinations of isotopes of the Er-Li systems, together with their physical parameters. The dipole lengths are comparable to the Er-Er van der Waals length $a_{dd}/r_{\mathrm{vdw}}\simeq 0.87$, which lie outside the perturbative regime $a_{dd}/r_{\mathrm{vdw}}\lesssim 0.5$ discussed in Sec. III.2. Nevertheless, they are not strong enough to be in the strong-dipole limit discussed in Sec. III.4; they are in the parameter region well described by the renormalized van der Waals universality shown in Sec. III.3. We can therefore use Eq. (15) to evaluate the binding energy of the Efimov states for the bosonic Er atoms interacting with a 6Li atom in unitary limit. The $s$-wave scattering length between the Er atoms $a^{(\mathrm{HH})}$ generally depends on the applied magnetic field, owing to the existence of a series of broad and narrow Feshbach resonances [57, 65], and it is yet to be determined by experiments or by sophisticated theoretical calculations. Still, by using the background $s$-wave scattering lengths reported in the literature [57, 58] as representative values, we can estimate the binding energy of the Efimov states $|E|=\kappa_{*}^{2}/M$. The results are tabulated in Table 1; we find that the difference in the values of $a^{(\mathrm{HH})}$ significantly affects the three-body parameter $\kappa_{*}$ across the isotopes while the mass and $a_{dd}$ are almost the same and therefore play marginal roles. The values of $\kappa_{*}$ in Table 1 can be readily improved with better knowledge of the Er-Er scattering length $a^{(\mathrm{HH})}$. We also note that the larger binding energy states in Table 1 may not be so accurate because we use the low-energy analytical formula in Eq. (15); this can also be straightforwardly improved using the numerical results of Sec. III.3.
Table 1: Three-body parameters of the Er-Er-Li Efimov states: binding energy at the unitary Er-Li scattering length $|E|=\kappa_{*}^{2}/M$, and the Er-Li scattering length $a_{-}^{(\mathrm{HL})}$ ($a_{*}^{(\mathrm{HL})}$) at which the Efimov states dissociate into three atoms (an atom and a dimer), respectively. The van der Waals and dipole lengths between the Er atoms are from Refs [57, 58] with the mass scaling. $a^{(\mathrm{HH})}$ between the heavy atoms are estimated by the background scattering lengths $a_{bg}$ tabulated in Refs. [58]. $a_{0}$ is the Bohr radius.
@@ V Conclusion & Outlook
We have studied the universality of the Efimov states in a mass-imbalanced three-body system of heavy-heavy-light atoms with a unitary interaction between the heavy-light atoms, with the major motivation of elucidating the interplay between the van der Waals and magnetic dipole interactions in Er-Er-Li Efimov states. By reducing the three-body problem to a coupled-channel two-body equation between the heavy atoms using the Born-Oppenheimer approximation, we have investigated analytically and numerically the three-body parameters (i.e., binding energies in the unitary limit) of the Efimov states for various strengths of dipole interactions between the bosonic and fermionic heavy atoms.
@@
In the absence of dipole interaction, using the quantum defect theory, we derive analytical formulas of the three-body parameters of the Efimov states, which are represented as explicit universal functions of the van der Waals length and $s$-wave scattering length ($p$-wave scattering volume) between the heavy bosonic (fermionic) atoms. The analytical formulas excellently reproduce the numerical results reported in Ref. [25], demonstrating the van der Waals universality in an analytical manner.
@@
In the presence of dipole interaction, different orbital angular momentum states get coupled, thus we have performed coupled-channel numerical calculations. For a weak dipole interaction, we find that the three-body parameters are affected by the dipole interaction in a different manner for the bosons and fermions. Due to the presence (absence) of the dipole matrix element for bosons in the $L=0$ (fermions in $L=1$) states, the three-body parameter gets a quadratic (linear) shift with respect to a perturbative dipole interaction for the bosons (fermions), respectively.
@@
When the dipole interaction is as strong as the van der Waals interaction, we find that the values of the three-body parameters get significantly altered from those of a non-dipolar system. However, once the $s$-wave scattering length between the heavy atoms is calculated in the presence of both the dipole and van der Waals interactions, the QDT formulas derived for a pure van der Waals system can well reproduce the three-body parameters of the bosonic Efimov states. Using this renormalized van der Waals universality, we have estimated the three-body parameters of the Er-Er-Li Efimov states for the bososnic Er isotopes. For the fermionic heavy atoms, on the other hand, the $p$-wave scattering volume can no longer be defined, owing to the long-range nature of the dipole interaction, so that the renormalized van der Waals universality does not hold. However, they can still be described universally with a low-energy scattering parameter between the two fermionic dipoles which characterizes the phase shift of an asymptotic $p$-wave state, so that the universality of the three-body parameter holds true.
@@
For an extremely large dipole interaction, neither the $s$-wave scattering length nor the $p$-wave low-energy scattering parameter can universally characterize the Efimov states, because the two dipoles get preferentially aligned in a linear geometry at short distance and they are neither $s$-wave nor $p$-wave dominant. In this limit, a quasi-one-dimensional scattering parameter is capable of capturing the short-range scattering phase. We show that except for some regions between avoided crossings, the Efimov states can be universally described by the one-dimensional scattering parameter.
@@
With the above studies, we have elucidated how the interplay of the van der Waals and dipole interactions can affect the universality of the Efimov states. In particular, we have clarified the roles of the statistics of the particles and the angular momentum channels, in the presence of both isotropic and anisotropic forces. Our work is not only relevant for the current challenge of observing Efimov states in the highly mass-imbalanced cold atom mixture of Er-Li atoms [41, 42], but also can shed new light on halo nuclear phenomena. The Efimov states are believed to appear universally in various strongly interacting quantum systems [18, 8, 3]. In particular, Efimov-like weakly bound three-body states have been actively studied in nuclear systems, such as neutron-rich nuclei [7, 80, 81, 82], the Hoyle state of 12[83, 84], and in near-breakup excited states of nuclei [85]. In nuclear physics, the tensor force has been pointed out to play a crucial role in determining the stability and reactions of nuclei. Notably, the tensor force has an anisotropic nature similar to the dipole-dipole interaction [86]. Investigating the interplay of the van der Waals interaction and dipole interaction is thus analogous to the nuclear problems where the interplay of the central part of the strong force and the tensor force is relevant. Our study of the dipole and van der Waals interactions in strongly interacting quantum few-body systems lays the basis for the quantum simulation of nuclear phenomena using dipolar cold-atom mixtures [41, 42, 43, 58, 87].